OpenStudy (anonymous):

1. There are 6 numbers on a normal die. What is the probability that if you roll the die, you'll get a 5 or a 6? Express your answer as a fraction reduced to its simplest form. 2. There are six numbers on a normal die. What is the probability that if you roll the die, you'll get a an even number? Express your answer as a fraction reduced to its simplest form. Just a little quick help please! :)

OpenStudy (anonymous):

A) sample space = {1,2,3,4,5,6}, wanted: {5} or {6} = 1/6+1/6=1/3 B) SS = {1,2,3,4,5,6} wanted: {2,4,6} so 3/6=1/2

OpenStudy (anonymous):

can you explain A/1. a little better for me please?

OpenStudy (anonymous):

Ok, we have 6 possibilities right ? It can be 1 .. to 6; but the question says that you get a "5" or a "6", now let's think about this: "or" is a keyword which tells you to multiply each outcome because they are independent events so we should first find the probability of having a "5" over 6 numbers, which is 1/6 and having a "6" over 6 numbers, which is 1/6, too, and then multiply them: 1/6+1/6=2/6=1/3

OpenStudy (anonymous):

ah, that makes sense. i just wanted to be clear for future situations. thank you so much!

OpenStudy (anonymous):

sorry add them i meant

OpenStudy (anonymous):

please change all "multiply"s to "add"s :)